OpenStudy (mimijola):

If XZ=50, what are XY and YZ XY=a YZ=a+8 Can someone help me ASAP..IM SO CONFUSED

OpenStudy (mimijola):

HELPPPPPPPPPPPP!!!!!!!!!!!!!

OpenStudy (mimijola):

PLEASE!

OpenStudy (mhchen):

Is there a diagram?

OpenStudy (mimijola):



OpenStudy (mhchen):

That's a LOT better. So XZ = XY + YZ right? You are given: XZ = 50 XY = a YZ = a + 8 So if you plug them into our equation: 50 = (a) + (a + 8) <--understand? 50 = a + a + 8 <-- Since it's addition, parentheses don't matter (Associative) 50 - 8 = a + a + 8 - 8 <-- subtracting 8 to BOTH sides, so it's still equal. 42 = a + a <-- simplified 42 = 2a <-- combining 'like-terms' 42/2 = 2a/2 <-- dividing BOTH sides by 2, so it's equal 21 = a <-- simplifying. So now that you know a = 21 XY = a YZ = a + 8 Can you figure it out now?

OpenStudy (mimijola):

Yes!!! Thank you so much!!!!